rectangle rug perimeter is 460ft the length is 9 times as long as the width what is the width and length
Answer by amarjeeth123(569)

Let the width be x
Then the length is 9x.
Rectangle perimeter=2(x+9x)=2(10x)=20x
20x=460
x=23 ft
The width is 23 ft.
The length is 207 ft.